OpenStudy (anonymous):

Factor the trinomial below. 7x2 - 42x - 49

OpenStudy (callisto):

7x^2 - 42x - 49 = 7 (x^2 -6x -7) = 7 (x^2 - 7x +x-7) = 7 [ x(x-7) + (x-7)] = 7 (x+1)(x-7)